OpenStudy (anonymous):

I need help understanding functions. Determine whether function is one-to-one and graph: f(x)=-2x-7

OpenStudy (anonymous):

 my teacher solved like this, and said that, Yes it is 1-1. explain?

OpenStudy (anonymous):

I guess I'm just having touble remembering how to solve a function

OpenStudy (anonymous):

1-1 meand that wherever you cut through the graph of the function, you'll get only one solution.

OpenStudy (anonymous):

 For the above curve, you'll get two solutions when you cut through it.

OpenStudy (anonymous):

*means

OpenStudy (anonymous):

So the function of this curve is not 1-1.

OpenStudy (anonymous):

where do the points (0,-7) and (-9,2) come from?

OpenStudy (anonymous):

I get that -2 an -7 are the -9, but what about the rest?

OpenStudy (anonymous):

I am not sure but according o the graph you drew, the points I see are (0,-7) and (2,-9). The first one is correct beacuse when you replace 0 in the function i.e f(x)=-2x-7 y=-2(0)-7 y=-7 So when x=0, y=-7. There seems to be a mistake with your second set of points. Because if you replace x=2 in the function, you get y=-2(2)-7 y=-4-7 y=-11 --> (2,-11).